OpenStudy (anonymous):

Find the derivative of 4arcsin(x/3)

OpenStudy (anonymous):

The derivative of y = asinx is y’ = 1/sqrt(1-x^2) . Therefore, y = 4 asin(x^3), y’ =4•[ 1/sqrt(1 – (x^3)^2)] •3x^2 = 12x^2 / sqrt(1 - x^6)